### v2

@@ Appendix A $A^{C}_{v}$ for different ship types
Table 3 collects these averages. In line with our intuition, cargo ships’ routes cover the largest average area of all ship types, as one of their primary goals is to ship goods effectively over long distances. Interestingly, Tug and tow boats cover the 3rd highest convex hull area. This finding might initially seem paradoxical given their role in assistive navigation in ports or terminals, typically associated with short and compressed trajectories. However, this observation can be rationalized by considering that our analysis only includes trajectories with more than 50 messages and a minimum speed of $1kn$. A visual inspection using heatmaps in Figures 14 and 15 revealed that Tug and Towboats travel longer routes than expected under these circumstances. On the other hand, High-Speed-Crafts (HSC) are mostly used as high-speed ferries for short passages in commercial contexts, which explains their relatively small area covered.
@@
Appendix B Quantile comparison results
@@ Appendix C Python package implementation details
{minted}
[linenos,frame=lines,fontsize=]python from pytsa import decode
decode( source = "path/to/raw_dir", dest = "path/to/decoded_dir", njobs = 1 )
@@
If, instead, it is desired for the data to undergo a pre-processing step, a pre-processing function that all messages must pass through before extraction can be defined. For example, to apply the speed constraint from Section 3.2, the instantiation can be performed as described in Code Listing C.1.
{minted}
[linenos,frame=lines,fontsize=]python import pandas as pd import pytsa
from pathlib import Path
# Column name for SOG in the decoded # csv data. _speed = …
# Filter out all vessels that are slower than 1 knot # and faster than 30 knots. # Any pre-processor is mandated to take in a single # pandas DataFrame and return a single pandas DataFrame def speed_filter(df: pd.DataFrame) -> pd.DataFrame: return df[(df[_speed] > 1) & (df[_speed] < 30)]
# Bounding Box for the area of study in this research frame = pytsa.BoundingBox( LATMIN = 52.2, # [°N] LATMAX = 56.9, # [°N] LONMIN = 6.3, # [°E] LONMAX = 9.5, # [°E] ) # This can be either a single file or a # list of files dynamic_data = Path("/path/to/dynamic.csv") static_data = Path("/path/to/static.csv")
search_agent = pytsa.SearchAgent( msg12318file = dynamic_data, msg5file = static_data frame = frame, preprocessor = speed_filter # Apply the speed filter using the "preprocessor" keyword )
After creating the instance, the extract_all() method can be used to construct and, if desired, also split the trajectories according to the split-point procedure from Section 4.2. The method returns a Python dictionary with every unique MMSI as keys and their corresponding TargetShip objects as values. These objects contain information about the ship type and length and all trajectories identified as belonging to this MMSI. The method has an integrated switch to turn the split-point procedure on or off. If turned off, every returned TargetShip only contains a single trajectory, consisting of all AIS messages sent from its MMSI sorted by time.
We look at the code used to produce Figure 11 to demonstrate this approach’s capabilities. We also utilize the visualization module shipped with PyTSA for this approach. {minted}[linenos,frame=lines,fontsize=,mathescape]python # # This listing continues Code Listing C.1 # # To perform only trajectory construction without # using the split-point procedure, we use the # ‘skip_tsplit=True‘ option. ships = search_agent.extract_all(njobs=16,skip_tsplit=True)
# Define the area around Aabenraa AABENRAA = pytsa.BoundingBox( LATMIN = 54.9, # [°N] LATMAX = 55.5, # [°N] LONMIN = 9.2, # [°E] LONMAX = 10.0, # [°E] )
# Plot trajectories on the map for the area # around AABENRAA plot_trajectories_on_map( ships = ships, extent = AABENRAA )
# With split-point procedure applied ships = search_agent.extract_all(njobs=4,skip_tsplit=False) plot_trajectories_on_map( ships = ships, extent = AABENRAA )
@@
# Construct a Recipe object to assess every trajectory based on the rule functions # given to it. In this case, we use the partial class to coerce the rules to only take a single argument, # which is the trajectory given to it. assessment = Recipe( partial(too_few_obs, n = 50), partial(convex_hull_area, area = 3e5) )
# Instantiate the Inspector and split the trajectories into accepted and rejected ones. # In this example, the rejected trajectories are discarded. accepted, rejected = pytsa.Inspector(data = ships, recipe = assessment).inspect()
# Save a 500x500 px heatmap for every ship type types = [t for t in pytsa.ShipType] names = [t.name for t in types] expanded = [] for t in types: if isinstance(t.value, int): expanded.append([t.value]) else: expanded.append(list(t.value)) for i,t in enumerate(expanded): a = mmsi:s for mmsi,s in accepted.items() if any(st in t for st in s.ship_type) binned_heatmap(targets = a, bb = frame, npixels = 500, title = f"Heatmap for names[i]")

### v3

@@ Appendix A $A^{C}_{v}$ for different ship types
Table 8 collects these averages. In line with our intuition, cargo ships’ routes cover the largest average area of all ship types, as one of their primary goals is to ship goods effectively over long distances. Interestingly, Tug and tow boats cover the 3rd highest convex hull area. This finding might initially seem paradoxical given their role in assistive navigation in ports or terminals, typically associated with short and compressed trajectories. However, this observation can be rationalized by considering that our analysis only includes trajectories with more than 50 messages and a minimum speed of $1kn$. A visual inspection using heatmaps in Figures 16 and 17 revealed that Tug and Towboats travel longer routes than expected under these circumstances. On the other hand, High-Speed-Crafts (HSC) are primarily used as high-speed ferries for short passages in commercial contexts, which explains their relatively small area covered.
@@
Appendix B Time scale of empirical distributions
The split-point methodology outlined in Section 6.2 requires the computation of five distinct empirical distributions, using their quantiles to establish thresholds. In this process, we explicitly used trajectory data from one year of AIS records as the basis for each distribution. To investigate how much data is needed to accurately calculate the empirical distribution functions 18 displays the (ship-length-independent) quantile functions for all metrics, each calculated using different amounts of trajectory data (given as days and months). Visually, these distributions appear somewhat similar, leading to the assumption that the data follows identical distributions over various time horizons.
The computed quantile values in Table 9 only show minor differences in absolute terms, which leaves leeway to the choice of the amount of data used to calculate the quantiles. However, due to the huge sample sizes (over $10^{6}$ observations) that lead to very narrow confidence intervals, almost all quantiles are tested statistically differently. Ultimately, we use one year of data throughout the study to maximize generalizability.
@@ Appendix C Python package implementation details
1 import pandas as pd
2 import pytsa
3
4 from pathlib import Path
5
6 # Column name for SOG in the decoded
7 # csv data.
8 _speed =
9
10 # Filter out all vessels that are slower than 1 knot
11 # and faster than 30 knots.
12 # Any pre-processor is mandated to take in a single
13 # pandas DataFrame and return a single pandas DataFrame
14 def speed_filter(df: pd.DataFrame) -> pd.DataFrame:
15 return df[(df[_speed] > 1) & (df[_speed] < 30)]
16
17 # Bounding Box for the area of study in this research
18 frame = pytsa.BoundingBox(
19 LATMIN = 52.2, # [°N]
20 LATMAX = 56.9, # [°N]
21 LONMIN = 6.3, # [°E]
22 LONMAX = 9.5, # [°E]
23 )
24 # This can be either a single file or a
25 # list of files
26 dynamic_data = Path("/path/to/dynamic.csv")
27 static_data = Path("/path/to/static.csv")
28
29 search_agent = pytsa.SearchAgent(
30 msg12318file = dynamic_data,
31 msg5file = static_data
32 frame = frame,
33 preprocessor = speed_filter # Apply the speed filter using the "preprocessor" keyword
34 )
@@
1from pytsa import decode
2
3decode(
4 source = "path/to/raw_dir",
5 dest = "path/to/decoded_dir",
6 njobs = 1
7)
@@
If, instead, it is desired for the data to undergo a pre-processing step, a pre-processing function that all messages must pass through before extraction can be defined. For example, to apply the speed constraint from Section 5.2, the instantiation can be performed as described in Code Listing C.1. After creating the instance, the extract_all() method can be used to construct and, if desired, also split the trajectories according to the split-point procedure from Section 6.2. The method returns a Python dictionary with every unique MMSI as keys and their corresponding TargetShip objects as values. These objects contain information about the ship type and length and all trajectories identified as belonging to this MMSI. The method has an integrated switch to turn the split-point procedure on or off. If turned off, every returned TargetShip only contains a single trajectory, consisting of all AIS messages sent from its MMSI sorted by time.
We look at the code used to produce Figure 12 to demonstrate this approach’s capabilities. We also utilize the visualization module shipped with PyTSA for this approach.
1#
[…]
3#
4# To perform only trajectory construction without
5# using the split-point procedure, we use the
6# ‘skip_tsplit=True‘ option.
7ships = search_agent.extract_all(njobs=16,skip_tsplit=True)
8
9# Define the area around Aabenraa
10AABENRAA = pytsa.BoundingBox(
11 LATMIN = 54.9, # [°N]
12 LATMAX = 55.5, # [°N]
13 LONMIN = 9.2, # [°E]
14 LONMAX = 10.0, # [°E]
15)
16
17# Plot trajectories on the map for the area
18# around AABENRAA
19plot_trajectories_on_map(
20 ships = ships,
21 extent = AABENRAA
22)
23
24# With split-point procedure applied
25ships = search_agent.extract_all(njobs=4,skip_tsplit=False)
26plot_trajectories_on_map(
27 ships = ships,
28 extent = AABENRAA
29)
@@
With the help of the SearchAgent class, we could easily construct and split trajectories from raw AIS messages. The spatial assessment from Section 9 is implemented into PyTSA via the Inspector class, which filters the extracted trajectories based on a set of user-defined Rules. All Rules must have a function signature of def my_rule(track: Track) -> bool, where Track = list[AISMessage] is a list of AIS message objects defined in pytsa.structs. Rules must be defined such that they take in a single trajectory and return True if the trajectory is to be rejected. Rules must be combined to create a recipe given to the inspector for trajectory assessment. To showcase the functionality of the Inspector class and its usage with pre-defined Rules, we reconstruct the heatmaps from Figures 16 and 17 in the following code example.
1#
@@
3#
4# Construct and split trajectories in parallel
5ships = search_agent.extract_all(njobs=4)
6
7# Construct a Recipe object to assess every trajectory based on the rule functions
8# given to it. In this case, we use the partial class to coerce the rules to only take a single argument,
9# which is the trajectory given to it.
10assessment = Recipe(
11 partial(too_few_obs, n = 50),
12 partial(convex_hull_area, area = 3e5)
13)
14
15# Instantiate the Inspector and split the trajectories into accepted and rejected ones.
16# In this example, the rejected trajectories are discarded.
17accepted, rejected = pytsa.Inspector(data = ships, recipe = assessment).inspect()
18
19# Save a 500x500 px heatmap for every ship type
20types = [t for t in pytsa.ShipType]
21names = [t.name for t in types]
22expanded = []
23for t in types:
24 if isinstance(t.value, int):
25 expanded.append([t.value])
26 else:
27 expanded.append(list(t.value))
28for i,t in enumerate(expanded):
29 a = {mmsi:s for mmsi,s in accepted.items() if any(st in t for st in s.ship_type)}
30 binned_heatmap(targets = a, bb = frame, npixels = 500, title = f"Heatmap for {names[i]}")